# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 3.3 Finite Unions of CSPs
The index of $\sim$ is smaller than or equal to the degree of $\Phi$.
Lemma 3.17.
If the complement of $\llbracket\Phi\rrbracket$ is closed under homomorphisms, then there are finitely many GSO $\tau$-sentences $\Phi_{1},\dots,\Phi_{t}$ each of which describes a CSP such that $\Phi$ is equivalent to $\Phi_{1}\vee\cdots\vee\Phi_{t}$. If $\Phi$ is an MSO sentence, then $\Phi_{1},\dots,\Phi_{t}$ can be chosen to be MSO sentences as well.
Proof 3.18.
We prove the statement by induction on the degree $n$ of $\Phi$. By Lemma 3.16 the equivalence relation $\sim$ has at most $n$ equivalence classes on $\tau$-structures. Hence, if $n=1$, then $\llbracket\Phi\rrbracket$ is closed under disjoint unions, and we are done.
Let $\mathfrak{A}_{1},\dots,\mathfrak{A}_{s}$ be $\tau$-structures such that $\{[\mathfrak{A}_{1}],\dots,[\mathfrak{A}_{s}]\}$ is the set of all equivalence classes of $\sim$ that are distinct from $[\mathfrak{I}]$. Let $\mathcal{D}_{i}$ be the smallest subclass of $\llbracket\Phi\rrbracket$ that contains $[\mathfrak{A}_{i}]$ and whose complement is closed under homomorphisms. Note that $\llbracket\Phi\rrbracket=\bigcup_{i\leq s}{\mathcal{D}}_{i}$ since $[\mathfrak{I}]$ is contained in ${\mathcal{D}}_{i}$ for all $i\leq s$. By Lemma 3.11 (1), each $\mathcal{D}_{i}$ is a union of $\sim$-classes which are themselves a union of $\equiv_{q}^{\operatorname{GSO}}$-classes by Lemma 3.14. It follows that there exists $I_{i}\subseteq\{1,\ldots,m\}$ such that ${\mathcal{D}}_{i}=\bigcup_{j\in I_{i}}\llbracket\Psi_{j}\rrbracket$. We define $\Phi_{i}:=\bigvee_{j\in I_{i}}\Psi_{j}$. Note that the GSO sentence $\Phi_{i}$ is of quantifier rank $q$ such that ${\mathcal{D}}_{i}=\llbracket\Phi_{i}\rrbracket$. Hence, $\Phi$ is equivalent to $\bigvee_{i\leq s}\Phi_{i}$. Lemma 3.11 (2) asserts that $\llbracket\Phi\rrbracket\setminus{\mathcal{D}}_{i}$ is non-empty, and hence the degree of $\Phi_{i}$ must be strictly smaller than $n$ for all $i\in\{1,\ldots,s\}$. The statement now follows from the inductive assumption. The same argument applies to MSO as well.
Lemma 3.17 together with Corollary 3.8 implies the following.
@@
Every GSO sentence which is closed under homomorphisms is equivalent to a finite conjunction of GSO sentences each of which describes the complement of a CSP of a countable $\omega$-categorical structure. The analogous statement holds for MSO.
Not every homomorphism-closed class of structures that can be expressed in Second-order Logic is a finite intersection of complements of CSPs. We even have an example of a class of finite $\tau$-structures that can be expressed in Datalog but cannot be written in this form.
@@
It is easy to see that each of the structures in $\{\mathfrak{P}_{n}\mid n\geq 1\}$ is not contained in $\mathcal{C}$, and that the disjoint union of $\mathfrak{P}_{i}$ and $\mathfrak{P}_{j}$, for $i\neq j$, is contained in ${\mathcal{C}}$. It follows that $\mathcal{C}$ is not a finite intersection of complements of CSPs (and, by Corollary 3.19, cannot be expressed in GSO).
@@ 4 Canonical Datalog Programs
A remarkable fact about the expressive power of Datalog for constraint satisfaction problems over finite domains is the existence of canonical Datalog programs and the characterisation of Datalog expressivity in terms of an existential pebble game [29]; this has been generalised to CSPs for $\omega$-categorical structures.
4.1 The Canonical Program of Width $(\ell,k)$
The following concept is due to Feder and Vardi [29] for complements of finite-domain CSPs and has been generalised to complements of $\omega$-categorical CSPs by Bodirsky and Dalmau [12].
@@ 4.1 The Canonical Program of Width $(\ell,k)$
Let $\mathfrak{B}$ be an $\omega$-categorical structure with a finite relational signature $\tau$. Let $\mathfrak{B}^{\prime}$ be the expansion of $\mathfrak{B}$ by all primitively positively definable relations of arity at most $\ell$ and let $\tau^{\prime}$ be the (finite) signature of $\mathfrak{B}^{\prime}$. Then the canonical $(\ell,k)$-Datalog program for $\mathfrak{B}$ has IDBs $\tau^{\prime}\setminus\tau$ and EDBs $\tau$. The empty 00-ary relation serves as ${\sf goal}$. There is a finite number of inequivalent formulas $\psi(\bar{x},\bar{y})$ of the form
@@
having at most $k$ variables and where $\psi_{1},\dots,\psi_{j}$ are atomic $(\tau\cup\tau^{\prime})$-formulas and $R\in\tau$. For each of the inequivalent formulas $\psi(\bar{x},\bar{y})$ such that $\mathfrak{B}^{\prime}\models\forall\bar{x}\exists\bar{y}.\psi(\bar{x},\bar{y})$ we introduce a rule
@@
Let $\mathfrak{B}$ be a countable $\omega$-categorical $\tau$-structure. Let $\ell,k\in{\mathbb{N}}$ and let $\Pi$ be the canonical Datalog program of width $(\ell,k)$ for the complement of $\operatorname{CSP}(\mathfrak{B})$. Then for every finite $\tau$-structure $\mathfrak{A}$ the following are equivalent:
@@
Moreover, $\Pi$ is also a canonical Datalog program of width $(\ell,k)$ in the sense defined in the introduction.
These results can be applied to finite unions of complements of CSPs for $\omega$-categorical structures because of the following well-known fact.
@@
We say that a Datalog rule $\psi_{0}{:}{-}\psi_{1},\dots,\psi_{m}$ is connected if $\{\psi_{1},\dots,\psi_{m}\}$ cannot be partitioned into two non-empty subsets with disjoint sets of variables.
@@
As we have mentioned in the introduction, it suffices to show that the complement of $\llbracket\Pi\rrbracket$ is closed under disjoint unions. So let $\mathfrak{A}$ and $\mathfrak{B}$ be structures where $\Pi$ does not derive the goal predicate. Let $\mathfrak{A}^{\prime}$ and $\mathfrak{B}^{\prime}$ be the structure computed by $\Pi$ on $\mathfrak{A}$ and on $\mathfrak{B}$, respectively. If the body of a rule of $\Pi$ holds on $\mathfrak{A}^{\prime}\uplus\mathfrak{B}^{\prime}$, then all variables must take values from $\mathfrak{A}^{\prime}$, or all variables must take values from $\mathfrak{B}^{\prime}$, because the rule is connected. Hence, $\Pi$ does not derive new facts on $\mathfrak{A}^{\prime}\uplus\mathfrak{B}^{\prime}$, and in particular does not derive the goal predicate. Hence $\mathfrak{A}\uplus\mathfrak{B}$ belongs to the complement of $\llbracket\Pi\rrbracket$ we well.
@@
Let $\mathfrak{B}$ be a countable $\omega$-categorical structure such that there exists a Datalog program $\Pi$ for the complement of $\operatorname{CSP}(\mathfrak{B})$. Then there also exists a connected Datalog program for the complement of $\operatorname{CSP}(\mathfrak{B})$. In fact, if $\Pi$ has width $(\ell,k)$, then such a program can be obtained from the canonical Datalog program of width $(\ell,k)$ by removing all rules that are not connected.
@@
Let $\Pi$ be the canonical Datalog program for $\operatorname{CSP}(\mathfrak{B})$ (Theorem 4.2). Suppose that $\Pi$ contains a rule which is not connected, e.g., a rule of the form $\psi_{0}{:}{-}\psi_{1},\psi_{2}$ where $\psi_{1}$ and $\psi_{2}$ are conjunctions of atomic formulas with disjoint sets of variables. We claim that then $\psi_{0}{:}{-}\psi_{1}$ or $\psi_{0}{:}{-}\psi_{2}$ must be a rule of $\Pi$ as well. Otherwise, for each $i\in\{1,2\}$ there would be a substructure of $\mathfrak{A}_{i}$ of $\mathfrak{B}$ that satisfies $\psi_{i}\wedge\neg\psi_{0}$. Then the disjoint union of $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$ satisfies $\psi_{1}\wedge\psi_{2}\wedge\neg\psi_{0}$, a contradiction to the definition of the canonical Datalog program.
@@ 4.2 The Existential Pebble Game
An informal definition of the existential pebble game was given in the introduction; we now present a formal definition of the concept of a (positional) winning strategies for Duplicator in this game.
@@
A winning strategy for Duplicator for the existential $(\ell,k)$-pebble game on two relational $\tau$-structures $\mathfrak{A}$, $\mathfrak{B}$ is a non-empty set $\mathcal{H}$ of partial homomorphisms from $\mathfrak{A}$ to $\mathfrak{B}$ such that
@@
In all our arguments about the existential pebble game, we only need this definition; in particular, we do not need the presentation of the existential pebble game as we have presented it in the introduction.
@@
    there exist $\ell,k\in{\mathbb{N}}$ such that for all finite structures $\mathfrak{A}$, Spoiler wins the $(\ell,k)$-game for $\llbracket\Phi\rrbracket$ on $\mathfrak{A}$ if and only if $\mathfrak{A}\models\Phi$.
@@
First suppose that $\llbracket\Phi\rrbracket$ is in Datalog. That is, there exists $\ell,k\in{\mathbb{N}}$ and a Datalog program $\Pi$ of width $(\ell,k)$ such that $\llbracket\Phi\rrbracket=\llbracket\Pi\rrbracket$. Then clearly $\llbracket\Phi\rrbracket$ is closed under homomorphisms, and by Lemma 3.17, there are GSO sentences $\Phi_{1},\dots,\Phi_{m}$ such that $\Phi$ is equivalent to $\Phi_{1}\wedge\cdots\wedge\Phi_{m}$ and $\llbracket\Phi_{i}\rrbracket$ is the complement of a CSP, for each $i\in\{1,\dots,m\}$. Corollary 3.8 implies that there exists an $\omega$-categorical structure $\mathfrak{B}_{i}$ such that $\operatorname{CSP}(\mathfrak{B}_{i})=\llbracket\neg\Phi_{i}\rrbracket$.
Suppose that $\mathfrak{A}$ is a finite $\tau$-structure such that $\mathfrak{A}\models\Phi$. Then Spoiler wins the $(\ell,k)$-game as follows. Suppose that Duplicator plays the countable structure $\mathfrak{B}$ such that $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi\rrbracket=\emptyset$. Then $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi_{i}\rrbracket=\emptyset$ for some $i\in\{1,\dots,m\}$; otherwise, if there is a structure $\mathfrak{A}_{i}\in\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi_{i}\rrbracket$ for every $i\in\{1,\dots,m\}$, then the disjoint union of $\mathfrak{A}_{1},\dots,\mathfrak{A}_{m}$ satisfies $\Phi_{i}$ since $\Phi_{i}$ is closed under homomorphisms, and is in $\operatorname{CSP}(\mathfrak{B})$ since $\operatorname{CSP}(\mathfrak{B})$ is closed under disjoint unions; but this is in contradiction to our assumption that $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi\rrbracket=\emptyset$. Hence, $\operatorname{CSP}(\mathfrak{B})\subseteq\operatorname{CSP}(\mathfrak{B}_{i})$ and there is a homomorphism $h$ from $\mathfrak{B}$ to $\mathfrak{B}_{i}$ (see [12]). Note that $\Pi$ is sound for the complement of $\operatorname{CSP}(\mathfrak{B}_{i})$, and $\Pi$ derives ${\sf goal}$ on $\mathfrak{A}$, and thus Theorem 4.2 implies that Spoiler wins the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B}_{i})$. But since $\mathfrak{B}$ homomorphically maps to $\mathfrak{B}_{i}$, this implies that Spoiler wins the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B})$.
@@
For the converse implication, suppose that 1. and 2. hold. Since $\llbracket\Phi\rrbracket$ is closed under homomorphisms, Corollary 3.19 implies that there are GSO sentences $\Phi_{1},\dots,\Phi_{m}$ and $\omega$-categorical structures $\mathfrak{B}_{1},\dots,\mathfrak{B}_{m}$ such that $\Phi$ is equivalent to $\Phi_{1}\wedge\cdots\wedge\Phi_{m}$ and $\llbracket\neg\Phi_{i}\rrbracket=\operatorname{CSP}(\mathfrak{B}_{i})$. By Theorem 4.2, for every $i\in\{1,\dots,m\}$ there exists a canonical Datalog program $\Pi_{i}$ of width $(\ell,k)$ for $\llbracket\Phi_{i}\rrbracket$. Then Lemma 4.3 implies that there exists a Datalog program $\Pi$ such that $\llbracket\Pi\rrbracket=\llbracket\Pi_{1}\rrbracket\cap\cdots\cap\llbracket\Pi_{m}\rrbracket$. Since each $\Pi_{i}$ is sound for $\llbracket\Phi_{i}\rrbracket$, it follows that $\Pi$ is sound for $\llbracket\Phi\rrbracket$. Hence, it suffices to show that if $\mathfrak{A}$ is a finite $\tau$-structure such that $\mathfrak{A}\models\Phi$, then $\Pi$ derives ${\sf goal}$ on $\mathfrak{A}$. Since $\mathfrak{A}\models\Phi_{i}$ for all $i\in\{1,\dots,m\}$, the assumption implies that Spoiler wins the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B}_{i})$. By Theorem 4.2, it follows that $\Pi_{i}$ derives ${\sf goal}$ on $\mathfrak{A}$. Hence, $\Pi$ derives ${\sf goal}$ on $\mathfrak{A}$.
5 A coNP-complete CSP in MSO
In this section we show that the class of CSPs in MSO is (under complexity-theoretic assumptions) larger than the class of CSPs for reducts of finitely bounded structures (see Section 3.1). Let $\mathcal{T}=\{\mathfrak{T}_{2},\mathfrak{T}_{3},\dots\}$ be the set of Henson tournaments: the tournament $\mathfrak{T}_{n}$, for $n\geq 2$, has vertices $10,1,\dots,n+$ 1 and the following edges:
  • $(i,i+1)$ for $i\in\{0,\dots,n\}$;
  • $(0,n+1)$;
  • $(j,i)$ for $i+1<j$ and $(i,j)\neq(0,n+1)$.
Proposition 5.1.
$\operatorname{CSP}(\mathfrak{H})$ can be expressed in MSO.
Proof 5.2.
We have to find an MSO sentence that holds on a given digraph $(V;E)$ if and only if $(V;E)$ does not embed any of the tournaments from $\mathcal{T}$. We specify an MSO $\{X,E\}$-sentence $\Phi$, for a unary relation symbol $X$, that is true on a finite $\{X,E\}$-structure $\mathfrak{S}$ if and only if $(X^{\mathfrak{S}};E^{\mathfrak{S}})$ is isomorphic to $\mathfrak{T}_{n}$, for some $n\geq 2$. In $\phi$ we existentially quantify over
  • two vertices $s,t\in X$ (that stand for the vertex 00 and the vertex $n+1$ in $\mathfrak{T}_{n}$).
The formula $\Phi$ has the following conjuncts:
  1. 1.
    a first-order formula that states that $E$ defines a tournament on $X$;
  2. 2.
  3. 3.
  4. 4.
    $E(s,t)$, $E(s,a)$, $E(a,b)$, and $E(x,s)$ for all $x\in X\setminus\{a,t\}$;
  5. 5.
  6. 6.
We claim that the MSO sentence $\forall x\colon\neg E(x,x)\wedge\forall X\colon\neg\Phi$ holds on a finite digraph if and only if the digraph is loopless and does not embed $\mathfrak{T}_{n}$, for all $n\geq 3$. The forwards implication easily follows from the observation that if $(X;T)$ is isomorphic to $\mathfrak{T}_{n}$, for some $n\geq 2$, then $\phi$ holds; this is straightforward from the construction of $\Phi$ (and the explanations above given in brackets). Conversely, suppose that $\Phi$ holds. Then $(X;T)$ is a tournament. We construct an isomorphism $f$ from $(X;T)$ to $\mathfrak{T}_{|X|-1}$ as follows. Define $f(s)\coloneqq 0$, $f(a)\coloneqq 1$, and $f(b)=2$. Since $E(a,b)$, by item 5 there exists exactly one $a^{\prime}\in A$ such that $E(b,a^{\prime})$ and $E(a^{\prime},a)$. Define $f(a^{\prime})\coloneqq 3$. If $a^{\prime}=t$ then we have found an isomorphism with $\mathfrak{T}_{2}$. Otherwise, the partial map $f$ defined so far is an embedding into $\mathfrak{T}_{n}$ for some $n\geq 3$. Item 6 and $E(b,a^{\prime})$ imply that there exists exactly one $b^{\prime}\in B$ such that $E(a^{\prime},b^{\prime})$ and $E(b^{\prime},b)$, and we define $f(b^{\prime})\coloneqq 4$. Continuing in this manner, we eventually define $f$ on all of $X$ and find an isomorphism with $\mathfrak{T}_{|X|-1}$.
This shows that $\operatorname{CSP}(\mathfrak{H})$ cannot be expressed, unless NP = coNP, as $\operatorname{CSP}(\mathfrak{B})$ for some reduct of a finitely bounded structure and such CSPs are in NP. We do not know how to show this statement without complexity-theoretic assumptions, even if we just want to rule out that $\operatorname{CSP}(\mathfrak{H})$ can be expressed as $\operatorname{CSP}(\mathfrak{B})$ for some reduct of a finitely bounded homogeneous structure.

### v3

@@ 3.3 Finite Unions of CSPs
Let $\Phi$ and $\sim$ be as in Lemma 3.19. Then the index of $\sim$ is smaller than or equal to the degree of $\Phi$.
Theorem 3.22.
Let $\Phi$ be a GSO sentence. If the complement of $\llbracket\Phi\rrbracket$ is closed under homomorphisms, then there are finitely many GSO $\tau$-sentences $\Phi_{1},\dots,\Phi_{t}$ each of which describes a CSP such that $\Phi$ is equivalent to $\Phi_{1}\vee\cdots\vee\Phi_{t}$. If $\Phi$ is an MSO sentence, then $\Phi_{1},\dots,\Phi_{t}$ can be chosen to be MSO sentences as well.
Proof 3.23.
We prove the statement by induction on the degree $n$ of $\Phi$. By Corollary 3.21 the equivalence relation $\sim$ has at most $n$ equivalence classes on $\tau$-structures. Hence, if $n=1$, then $\llbracket\Phi\rrbracket$ is closed under disjoint unions, and we are done by Remark 1.2.
Let $\mathfrak{A}_{1},\dots,\mathfrak{A}_{s}$ be $\tau$-structures such that $\{[\mathfrak{A}_{1}],\dots,[\mathfrak{A}_{s}]\}$ is the set of all equivalence classes of $\sim$ that are distinct from $\mathfrak{I}$] and contained in $\llbracket\Phi\rrbracket$. Let $\mathcal{D}_{i}$ be the smallest subclass of $\llbracket\Phi\rrbracket$ that contains $\mathfrak{A}_{i}]$ and whose complement is closed under homomorphisms. Note that $\llbracket\Phi\rrbracket=\bigcup_{i\leq s}{\mathcal{D}}_{i}$ since $\mathfrak{I}$] is contained in ${\mathcal{D}}_{i}$ for all $i\leq s$. By Lemma 3.15 (1), each $\mathcal{D}_{i}$ is a union of $\sim$-classes which are themselves a union of $\equiv_{q}^{\operatorname{GSO}}$-classes by Lemma 3.19. It follows that there exists $I_{i}\subseteq\{1,\ldots,m\}$ such that ${\mathcal{D}}_{i}=\bigcup_{j\in I_{i}}\llbracket\Psi_{j}\rrbracket$. We define $\Phi_{i}:=\bigvee_{j\in I_{i}}\Psi_{j}$. Note that the GSO sentence $\Phi_{i}$ is of quantifier rank $q$ such that ${\mathcal{D}}_{i}=\llbracket\Phi_{i}\rrbracket$. Hence, $\Phi$ is equivalent to $\bigvee_{i\leq s}\Phi_{i}$. Lemma 3.15 (2) asserts that $\llbracket\Phi\rrbracket\setminus{\mathcal{D}}_{i}$ is non-empty, and hence the degree of $\Phi_{i}$ must be strictly smaller than $n$ for all $i\in\{1,\ldots,s\}$. The statement now follows from the inductive assumption. The same argument applies to MSO as well.
By taking complements, Theorem 3.22 together with Corollary 3.10 implies the following.
@@
Every GSO sentence whose finite models are closed under homomorphisms is equivalent to a finite conjunction of GSO sentences each of which describes the complement of a CSP of a countable $\omega$-categorical structure. The analogous statement holds for MSO.
At this point, it seems worthwhile to stress that this property of GSO does not extend to full SO: Not every homomorphism-closed class of structures that can be expressed in Second-order Logic is a finite intersection of complements of CSPs. We present an example of a class of finite $\tau$-structures that can even be expressed in Datalog but cannot be written in this form.
@@
It is easy to see that each of the structures in $\{\mathfrak{P}_{n}\mid n\geq 1\}$ is not contained in $\mathcal{C}$, and that the disjoint union of $\mathfrak{P}_{i}$ and $\mathfrak{P}_{j}$, for $i\neq j$, is contained in $\mathcal{C}$}. It follows that $\mathcal{C}$ is not a finite intersection of complements of CSPs (and, by Corollary 3.24, cannot be expressed in GSO). $\triangle$
@@ 4 Canonical Datalog Programs
A remarkable fact about the expressive power of Datalog for constraint satisfaction problems over finite domains is the existence of canonical Datalog programs and the characterisation of Datalog expressivity in terms of an existential pebble game [32]; this has been generalised to CSPs for $\omega$-categorical structures [14]. In this section, we build on these results to characterise Datalog expressibility of GSO queries.
4.1 The Canonical Datalog Program of Width $(\ell,k)$
A Datalog program $\Pi$ is called sound for a class of $\tau$-structures $\mathcal{C}$} if $\llbracket\Pi\rrbracket\subseteq{\mathcal{C}}$. It is called a canonical Datalog program of width $(\ell,k)$ for $\mathcal{C}$}, for $\ell,k\in{\mathbb{N}}$, if if
  • $\Pi$ is sound for $\mathcal{C}$}, and
The following definition is due to Feder and Vardi [32] for finite structures and has been generalised to $\omega$-categorical structures by Bodirsky and Dalmau [14].
@@ 4.1 The Canonical Datalog Program of Width $(\ell,k)$
Let $\mathfrak{B}$ be an $\omega$-categorical structure with a finite relational signature $\tau$. Let $\mathfrak{B}^{\prime}$ be the expansion of $\mathfrak{B}$ by all primitively positively definable relations of arity at most $\ell$ and let $\tau^{\prime}$ be the (finite) signature of $\mathfrak{B}^{\prime}$. Then the canonical $(\ell,k)$-Datalog program for $\mathfrak{B}$ has IDBs $\tau^{\prime}\setminus\tau$ and EDBs $\tau$. The empty 0-ary relation serves as $\sf goal$}. There is a finite number of inequivalent formulas $\psi(\bar{x},\bar{y})$ of the form
@@
having at most $k$ variables and where $\psi_{1},\dots,\psi_{j}$ are atomic $\tau^{\prime}$-formulas and $R\in\tau^{\prime}\setminus\tau$. For each of the inequivalent formulas $\psi(\bar{x},\bar{y})$ such that $\mathfrak{B}^{\prime}\models\forall\bar{x},\bar{y}.\psi(\bar{x},\bar{y})$ we introduce a rule
@@
Let $\mathfrak{B}$ be a countable $\omega$-categorical $\tau$-structure. Let $\ell,k\in{\mathbb{N}}$ and let $\Pi$ be the canonical Datalog program of width $(\ell,k)$ for $\mathfrak{B}$. Then for every finite $\tau$-structure $\mathfrak{A}$ the following are equivalent:
@@
Moreover, $\Pi$ is a canonical Datalog program of width $(\ell,k)$ for the complement of $\operatorname{CSP}(\mathfrak{B})$.
These results extend to finite unions of complements of CSPs for $\omega$-categorical structures (and therefore to all homomorphism-closed GSO-expressible classes of structures by Corollary 3.24) because of the following well-known fact.
@@
If a GSO sentence is in Datalog, then whether or not the problem describes a CSP can be characterised by a syntactic description of the Datalog program (Proposition 4.9). We say that a Datalog rule $\psi_{0}{:}{-}\psi_{1},\dots,\psi_{m}$ is connected if $\{\psi_{1},\dots,\psi_{m}\}$ cannot be partitioned into two non-empty subsets with disjoint sets of variables. The following holds for Datalog in general.
@@
By Remark 1.2, it suffices to show that the complement of $\llbracket\Pi\rrbracket$ is closed under disjoint unions. So let $\mathfrak{A}$ and $\mathfrak{B}$ be structures where $\Pi$ does not derive the goal predicate. Let $\mathfrak{A}^{\prime}$ and $\mathfrak{B}^{\prime}$ be the structure computed by $\Pi$ on $\mathfrak{A}$ and on $\mathfrak{B}$, respectively. If the body of a rule of $\Pi$ holds on $\mathfrak{A}^{\prime}\uplus\mathfrak{B}^{\prime}$, then all variables must take values from $\mathfrak{A}^{\prime}$, or all variables must take values from $\mathfrak{B}^{\prime}$, because the rule is connected. Hence, $\Pi$ does not derive new facts on $\mathfrak{A}^{\prime}\uplus\mathfrak{B}^{\prime}$, and in particular does not derive the goal predicate. Hence $\mathfrak{A}\uplus\mathfrak{B}$ belongs to the complement of $\llbracket\Pi\rrbracket$ we well.
In the specific situation that a Datalog program describes a problem in GSO, Proposition 4.7 has a converse.
@@
Suppose that $\Pi$ is a Datalog program such that $\llbracket\Pi\rrbracket$ is in GSO and equals the complement of a CSP. Then there exists a connected Datalog program $\Pi^{\prime}$ defining $\llbracket\Pi\rrbracket$. In fact, if $\mathfrak{B}$ is the $\omega$-categorical structure such that $\llbracket\Pi\rrbracket$ is the complement of $\operatorname{CSP}(\mathfrak{B})$ (which exists by Corollary 3.10), and if $\Pi$ has width $(\ell,k)$, then such a program $\Pi^{\prime}$ can be obtained from the canonical Datalog program of width $(\ell,k)$ for $\mathfrak{B}$ by removing all rules that are not connected.
@@
Let $\Pi$ be the canonical Datalog program of width $(\ell,k)$ for $\mathfrak{B}$ (Definition 4.1). Suppose that $\Pi$ contains a rule which is not connected, e.g., a rule of the form $\psi_{0}{:}{-}\psi_{1},\psi_{2}$ where $\psi_{1}$ and $\psi_{2}$ are conjunctions of atomic formulas with disjoint sets of variables. We claim that then $\psi_{0}{:}{-}\psi_{1}$ or $\psi_{0}{:}{-}\psi_{2}$ must be a rule of $\Pi$ as well. Otherwise, for each $i\in\{1,2\}$ there would be a substructure of $\mathfrak{A}_{i}$ of $\mathfrak{B}$ that satisfies $\psi_{i}\wedge\neg\psi_{0}$. Then the disjoint union of $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$ satisfies $\psi_{1}\wedge\psi_{2}\wedge\neg\psi_{0}$, a contradiction to the definition of the canonical Datalog program of $\mathfrak{B}$.
@@ 4.2 The Existential Pebble Game
The existence of canonical Datalog programs for GSO can be be shown via a characterisation of Datalog based on existential pebble games. An informal definition of the existential pebble game was given in the introduction; we now present a formal definition of the concept of (positional) winning strategies for Duplicator in this game.
@@
A winning strategy for Duplicator for the existential $(\ell,k)$-pebble game on a pair of relational $\tau$-structures $(\mathfrak{A},\mathfrak{B})$ is a non-empty set $\mathcal{H}$ of partial homomorphisms from $\mathfrak{A}$ to $\mathfrak{B}$ such that
@@
We say that Duplicator wins the existential $(\ell,k)$-pebble game for $\llbracket\Phi\rrbracket$ on $\mathfrak{A}$ if there exists a countable $\tau$-structure $\mathfrak{B}$ such that $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi\rrbracket=\emptyset$ and Duplicator has a winning strategy on $(\mathfrak{A},\mathfrak{B})$ as in Definition 4.11, and otherwise we say that Spoiler wins the game. In all our arguments about the existential pebble game, we only need this definition; in particular, we do not need the presentation of the existential pebble game as we have presented it in the introduction.
@@
    there exist $\ell,k\in{\mathbb{N}}$ such that for all finite structures $\mathfrak{A}$, Spoiler wins the existential $(\ell,k)$-pebble game for $\llbracket\Phi\rrbracket$ on $\mathfrak{A}$ if and only if $\mathfrak{A}\models\Phi$.
@@
If $\llbracket\Phi\rrbracket$ is in Datalog, then it is closed under homomorphisms. Hence, for both directions of the proof we may apply Corollary 3.24 to $\Phi$, which implies that there are GSO sentences $\Phi_{1},\dots,\Phi_{m}$ and $\omega$-categorical structures $\mathfrak{B}_{1},\dots,\mathfrak{B}_{m}$ such that $\Phi$ is equivalent to $\Phi_{1}\wedge\cdots\wedge\Phi_{m}$ and $\llbracket\neg\Phi_{i}\rrbracket=\operatorname{CSP}(\mathfrak{B}_{i})$. First suppose that $\llbracket\Phi\rrbracket$ is in Datalog. That is, there exist $\ell,k\in{\mathbb{N}}$ and a Datalog program $\Pi$ of width $(\ell,k)$ such that $\llbracket\Phi\rrbracket=\llbracket\Pi\rrbracket$. Suppose that $\mathfrak{A}$ is a finite $\tau$-structure such that $\mathfrak{A}\models\Phi$. Then Spoiler wins the $(\ell,k)$-game as follows. Suppose that $\mathfrak{B}$ is a countable structure such that $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi\rrbracket=\emptyset$. Then $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi_{i}\rrbracket=\emptyset$ for some $i\in\{1,\dots,m\}$; otherwise, if there is a structure $\mathfrak{A}_{i}\in\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi_{i}\rrbracket$ for every $i\in\{1,\dots,m\}$, then the disjoint union of $\mathfrak{A}_{1},\dots,\mathfrak{A}_{m}$ satisfies $\Phi_{i}$ since $\Phi_{i}$ is closed under homomorphisms, and is in $\operatorname{CSP}(\mathfrak{B})$ since $\operatorname{CSP}(\mathfrak{B})$ is closed under disjoint unions; but this is in contradiction to our assumption that $\operatorname{CSP}(\mathfrak{B})\cap\llbracket\Phi\rrbracket=\emptyset$. Hence, $\operatorname{CSP}(\mathfrak{B})\subseteq\operatorname{CSP}(\mathfrak{B}_{i})$ and there is a homomorphism $h$ from $\mathfrak{B}$ to $\mathfrak{B}_{i}$ (see [14, Lemma 2]). Note that $\Pi$ is sound for the complement of $\operatorname{CSP}(\mathfrak{B}_{i})$, and $\Pi$ derives $\sf goal$} on $\mathfrak{A}$, and thus Theorem 4.2 implies that Spoiler wins the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B}_{i})$. But since $\mathfrak{B}$ homomorphically maps to $\mathfrak{B}_{i}$, this implies that Spoiler wins the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B})$.
@@
For the converse implication, suppose that 1. and 2. hold. By Theorem 4.2, for every $i\in\{1,\dots,m\}$ there exists a canonical Datalog program $\Pi_{i}$ of width $(\ell,k)$ for $\llbracket\Phi_{i}\rrbracket$. Then Lemma 4.3 implies that there exists a Datalog program $\Pi$ such that $\llbracket\Pi\rrbracket=\llbracket\Pi_{1}\rrbracket\cap\cdots\cap\llbracket\Pi_{m}\rrbracket$. Since each $\Pi_{i}$ is sound for $\llbracket\Phi_{i}\rrbracket$, it follows that $\Pi$ is sound for $\llbracket\Phi\rrbracket$. Hence, it suffices to show that if $\mathfrak{A}$ is a finite $\tau$-structure such that $\mathfrak{A}\models\Phi$, then $\Pi$ derives $\sf goal$} on $\mathfrak{A}$. Since $\mathfrak{A}\models\Phi_{i}$ for all $i\in\{1,\dots,m\}$, the assumption implies that Spoiler wins the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{B}_{i})$. By Theorem 4.2, it follows that $\Pi_{i}$ derives $\sf goal$} on $\mathfrak{A}$. Hence, $\Pi$ derives $\sf goal$} on $\mathfrak{A}$.
An example application of Theorem 4.12 for showing that a certain MSO sentence cannot be expressed in Datalog can be found in the proof of Theorem 6.7.